trish30:

Find the missing value, a : {\left( {4\sqrt x + 3} \right)^2} = 16x + a\sqrt x + 9

Zarkon:

\[{\left( {4\sqrt x + 3} \right)^2} = 16x+24\sqrt{x}+9\]

Hero:

@trish30, when it comes to problems such as these, it is often useful to start with \(x = 0\).